OpenStudy (anonymous):

The US postal service delivered 7.14 x 1010 pieces of mail in the month of December and 3.21 x 1010 in the month of January. What is the total mail delivered for these two months?

OpenStudy (anonymous):

7.14 * 10^10 + 3.21 * 10^10

OpenStudy (anonymous):

= 10.35 * 10^10 = 1.035 * 10^11